# A coupled finite element-virtual element method for thermomechanical analysis of electronic packaging structures

### v1

@@ 1 Introduction
Numerical simulation methods have become essential for electronic packaging reliability analysis, offering cost-effective and reliable assessment capabilities. However, thermomechanical coupling analysis in electronic packaging faces significant computational challenges due to multi-scale geometric features and complex material interfaces in modern designs. Researchers have developed various numerical approaches to address these challenges, including the Finite Element Method (FEM) [ye2022practical], Boundary Element Method (BEM) [VALLEPUGAESPINOSA202028], Isogeometric BEM [gong2024application, yu2021thermal, gong2022isogeometric], FEM-BEM coupling [gong2025coupled, qin2022application], Phase Field scheme [gong2025phase, LONG2025110753], Scientific Machine Learning [gong2025pinn], Finite Difference Method [raszkowski2020analysis] and Finite Volume Method [khor2010fvm]. Recent advances have also introduced specialized approaches: Inamdar et al. [inamdar2024modelling] developed physics-based Digital Twin modeling for thermomechanical degradation prediction; Feng et al. [feng2023engineered] proposed smoothed finite element algorithms for electro-thermal-mechanical coupling; and Wang et al. [wang2023structural] integrated machine learning approaches for thermal stress reduction strategies.
@@
Despite FEM’s mature theoretical framework and high computational accuracy for regular structures with quality meshes, it faces significant challenges when analyzing multi-scale geometric configurations. Complex geometries require extensive mesh refinement and careful element partitioning to prevent distortion, while geometric scale variations necessitate transitional meshes that can generate millions of elements. This dramatically increases computational cost and may cause convergence issues. The Virtual Element Method (VEM) addresses these limitations through enhanced geometric flexibility [gong2025VEMthermomechanical]. VEM accommodates arbitrary polygonal elements that simplify complex geometry discretization and enables local refinement through non-conforming meshes without compromising overall mesh quality. This capability reduces element count and computational cost while preserving solution accuracy, making VEM particularly suitable for multi-scale problems where traditional meshing approaches become prohibitively expensive.
@@
The VEM was first introduced by Beirao da Veiga et al. [beirao2013basic] for Poisson’s equation and has since expanded across multiple engineering applications. Early validation studies by Artioli et al. [artioli2017arbitrary] demonstrated VEM’s accuracy through computational benchmarks including Cook’s membrane tests. Subsequent research [sorgente2022role, sorgente2022polyhedral, van2020virtual, van2021virtual] has confirmed VEM’s geometric robustness and convergence properties for arbitrary element shapes, including non-convex configurations. VEM applications have extended beyond linear elasticity to encompass hyperelastic materials [xu2024high, chi2017some], contact mechanics [shen20222, cihan2022virtual], elastoplasticity [hudobivnik2019low, wriggers2017low], and phase field fracture modeling [liu2023virtual]. Dhanush et al. [dhanush2019implementation] pioneered VEM implementation for thermomechanical coupling using Abaqus UEL framework. Recently, Gong et al. [gong2025VEMthermomechanical] analyzed thermoelastic problems using VEM, presenting both the standard virtual element method and a stabilization-free virtual element method [XU2024116826] for thermomechanical behavior analysis in electronic packaging structures with geometric multi-scale features. Their work demonstrated that VEM’s polygonal mesh flexibility enables localized mesh modifications without affecting global mesh structure, making it particularly effective for electronic packaging reliability analysis involving complex geometries. However, VEM requires projection operators and stabilization terms that increase algorithmic complexity and may result in higher local computational costs compared to conventional FEM.
@@ 2.1 Governing equations
$\nabla\cdot(\lambda\nabla T)=0,$ (3)
where $\lambda$ is the thermal conductivity and $T$ is the temperature field. The boundary conditions on $\partial\Omega=\partial\Omega_{T}\cup\partial\Omega_{q}$ are expressed as
@@
where $\mathbf{I}$ is the identity tensor, $\alpha$ is the coefficient of thermal expansion, and $T_{0}$ is the reference temperature.
@@ 2.2 Weak formulation
$a_{T}(T,\varphi)=\ell_{T}(\varphi),$ (12)
where $\varphi$ is the test function (virtual temperature variation). The corresponding bilinear and linear forms are defined as
@@ 3.1.1 Thermal analysis
For the finite element discretization of the thermal problem, we employ the weak formulation established in Sec. 2.2. The domain is discretized into finite elements such that $\Omega\equiv\text{int}\left(\bigcup_{i}E_{i}\right)$, where element $E_{i}$ is chosen as a four-node quadrilateral isoparametric element.
@@
The integrals are evaluated using $2\times 2$ Gaussian quadrature in the local coordinate system
$\mathbf{K}_{\mathrm{FE}}^{th,E}=\sum_{p=1}^{2}\sum_{q=1}^{2}w_{p}w_{q}\,\lambda(\xi_{p},\eta_{q})\,\mathbf{B}_{T}^{\mathrm{T}}(\xi_{p},\eta_{q})\mathbf{B}_{T}(\xi_{p},\eta_{q})\,\det(\mathbf{J}(\xi_{p},\eta_{q})),$ (23)
where $w_{p}$ and $w_{q}$ are the Gaussian weights ($w_{1}=w_{2}=1.0$), $(\xi_{p},\eta_{q})$ are the integration points ($\xi_{1}=\eta_{1}=-1/\sqrt{3}$, $\xi_{2}=\eta_{2}=1/\sqrt{3}$), and $\det(\mathbf{J})$ is the determinant of the Jacobian matrix.
@@ 3.1.2 Thermoelastic analysis
For the finite element discretization of the mechanical problem, we employ the weak formulation established in Sec. 2.2. Within each element, the displacement field is interpolated using the same shape functions as the thermal analysis
@@
The element mechanical stiffness matrix is computed using Gaussian quadrature
$\small\mathbf{K}_{\mathrm{FE}}^{u,E}=\int_{E}\mathbf{B}_{u}^{\mathrm{T}}\mathbb{D}\mathbf{B}_{u}\,\mathrm{d}E=\sum_{p=1}^{2}\sum_{q=1}^{2}w_{p}w_{q}\,\mathbf{B}_{u}^{\mathrm{T}}(\xi_{p},\eta_{q})\mathbb{D}\mathbf{B}_{u}(\xi_{p},\eta_{q})\,\det(\mathbf{J}(\xi_{p},\eta_{q})).$ (29)
@@ 3.2.1 Thermal analysis
In Eq. (43), the first term on the right-hand side represents the consistency term computed from the integral of projected functions in element $E$. The second term represents the stabilization term that accounts for the energy associated with the non-polynomial component of the solution.
@@
$\mathbf{K}_{E}^{th,\mathrm{s}}=\tau^{h}\,\mathrm{tr}\left(\mathbf{K}_{E}^{th,\mathrm{c}}\right)(\mathbf{I}-\bm{\Pi})^{T}(\mathbf{I}-\bm{\Pi}),$ (46)
@@ 3.2.2 Thermoelastic analysis
where $\mathbf{v}^{h}$ represents the discrete vector function within the element.
@@
where $\tilde{\bm{\Pi}}^{*\nabla}$ is the matrix formulation of the Ritz projection operator solved from Eq. (56). The thermal strain $\hat{\bm{\varepsilon}}_{\mathrm{th}}=\alpha(T-T_{0})[1,1,0]^{\mathrm{T}}$ is expressed in Voigt form.
@@ 4.3.2 Thermal analysis phase
The thermal analysis phase implements the heat conduction formulation established in Sec. 2.2. For the FE domain, standard shape functions are constructed and element thermal stiffness matrices $\mathbf{K}_{\mathrm{FE}}^{th,E}$ are computed using the procedures described in Sec. 3.1.1. Simultaneously, the VE domain employs projection operators to calculate element stiffness matrices $\mathbf{K}_{\mathrm{VE}}^{th,E}$ following the VEM formulation.
@@ 4.3.3 Thermoelastic analysis phase
The mechanical analysis phase utilizes the computed temperature field as thermal loading input, implementing the thermomechanical coupling described in Sec. 2.2. Following procedures analogous to the thermal analysis, shape functions are constructed and element mechanical stiffness matrices are computed for both domains: $\mathbf{K}_{\mathrm{FE}}^{u,E}$ for the finite element region and $\mathbf{K}_{\mathrm{VE}}^{u,E}$ for the virtual element region.
@@ 5.1 Square plate with circular hole
Three different numerical approaches were employed to solve this problem: VEM, FEM, and the proposed FE-VE coupling method. The domain is partitioned such that the regular regions employ finite elements while the area near the circular hole utilizes virtual elements to handle the geometric complexity. The resulting von Mises stress distribution contours are presented in Fig. 5. The stress contours demonstrate excellent agreement among all three methods, particularly around the critical stress concentration region near the circular hole. Fig. 5(a) shows the VEM solution with its characteristic polygonal mesh that naturally adapts to the curved boundary geometry. Fig. 5(b) presents the FEM solution employing a structured quadrilateral mesh, while Fig. 5(c) illustrates the FE-VE coupling approach that effectively combines the advantages of both methods. The stress magnitude distributions and overall stress patterns are highly consistent across all three approaches, confirming the accuracy and reliability of the proposed coupling methodology.
@@
The stress distributions along the line $L_{1}$ (given in Fig. 4) were extracted for FEM, VEM, and FE-VE method under different Dofs, as illustrated in Fig. 6. The results demonstrate excellent agreement among all three approaches, confirming the accuracy and consistency of the proposed coupling methodology. Additionally, the computational times for the three methods were compared, as presented in Table 2. From Table 2, it is evident that computational time increases with the number of degrees of freedom (nDofs) for all three methods, which is expected behavior. When comparing methods with similar Dof (approximately 10,700), VEM demonstrates slightly better computational efficiency than FEM, requiring 2.06 seconds compared to 2.4 seconds for FEM. This efficiency gain is primarily attributed to VEM’s ability to achieve comparable accuracy using fewer elements (2,595 vs 5,263), thereby reducing the computational overhead. The FE-VE coupling approach shows competitive performance, with computational times that fall between pure VEM and pure FEM solutions, while providing the geometric flexibility advantages of the hybrid approach.
@@
where $N$ represents the total number of nodes in the calculation, $\bm{\sigma}_{\text{num}}$ denotes the numerical solution of nodal stress, and $\bm{\sigma}_{\text{ref}}$ represents the reference solution of nodal stress.
@@
Fig. 7 presents the MRE convergence comparison for FEM, VEM, and FE-VE coupling method as the nDof increases from approximately 5,000 to 25,000. The convergence analysis reveals distinct performance characteristics for each method, with convergence rates indicated by the slope annotations in the figure. FEM demonstrates the fastest convergence with a rate of 0.43, showing the steepest error reduction from approximately $1.3\times 10^{-3}$ to $6.2\times 10^{-4}$. VEM exhibits a convergence rate of 0.338, with errors decreasing from $1.41\times 10^{-3}$ to $8.0\times 10^{-4}$, demonstrating consistent performance across the degree-of-freedom range. The FE-VE coupling method achieves a convergence rate of 0.442, reducing errors from $1.56\times 10^{-3}$ to $7.3\times 10^{-4}$, performing competitively with pure FEM. The convergence rates indicate that the FE-VE coupling method actually achieves slightly better convergence behavior than pure FEM, while VEM shows the most gradual convergence. All three methods demonstrate stable convergence trends, confirming the reliability of the proposed coupling strategy. The FE-VE approach maintains good accuracy while providing the additional advantage of geometric flexibility for complex domain geometries.
@@ 5.2 Thermoelastic analysis of thick-walled cylindrical structure
We analyze a homogeneous thick-walled cylinder under thermal loading to validate the FEM-VEM coupling method. The analytical solution for this problem serves as a benchmark for accuracy assessment. Fig. 8 illustrates the geometry and boundary conditions of this verification case.
@@
Figure 8: Thick-walled cylinder configuration: (a) Full geometry with inner radius $r_{a}$, outer radius $r_{b}$, and applied temperatures $T_{a}$ and $T_{b}$; (b) Mesh employed in the FEM-VEM coupled method
@@
where $N$ is the number of evaluation points, and $X_{e}$ and $X$ denote the analytical and numerical solutions, respectively.
@@
Fig. 9 presents the convergence analysis comparing FEM-VEM coupling, pure VEM, and pure FEM methods, showing RMS errors for temperature as functions of degrees of freedom. The convergence rates, indicated by the slope annotations in the figure, demonstrate that the VEM achieves a convergence rate of 1.02, while the FEM-VEM coupling method exhibits a rate of 1.01, both outperforming pure FEM with a rate of 0.92. The proposed FEM-VEM coupling method maintains nearly optimal convergence characteristics comparable to pure VEM, while providing the additional advantage of geometric flexibility through strategic domain partitioning. All three methods show consistent error reduction with mesh refinement, confirming the stability and reliability of the coupling approach.

### v2

@@ 1 Introduction
Numerical simulation methods have become essential for electronic packaging reliability analysis, offering cost-effective and reliable assessment capabilities. However, thermomechanical coupling analysis in electronic packaging faces significant computational challenges due to multi-scale geometric features and complex material interfaces in modern designs. Researchers have developed various numerical approaches to address these challenges, including the Finite Element Method (FEM) [ye2022practical], Boundary Element Method (BEM) [VALLEPUGAESPINOSA202028], Isogeometric BEM [yu2021thermal, gong2022isogeometric], FEM-BEM coupling [gong2025coupled, qin2022application], Phase Field scheme [gong2025phase, LONG2025110753], Scientific Machine Learning [gong2025pinn], Finite Difference Method [raszkowski2020analysis] and Finite Volume Method [khor2010fvm]. Recent advances have also introduced specialized approaches: Inamdar et al. [inamdar2024modelling] developed physics-based Digital Twin modeling for thermomechanical degradation prediction; Feng et al. [feng2023engineered] proposed smoothed finite element algorithms for electro-thermal-mechanical coupling; and Wang et al. [wang2023structural] integrated machine learning approaches for thermal stress reduction strategies.
@@
Despite FEM’s mature theoretical framework and high computational accuracy for regular structures with quality meshes, it faces significant challenges when analyzing multi-scale geometric configurations. Complex geometries require extensive mesh refinement and careful element partitioning to prevent distortion, while geometric scale variations necessitate transitional meshes that can generate millions of elements. This dramatically increases computational cost and may cause convergence issues. The Virtual Element Method (VEM) addresses these limitations through enhanced geometric flexibility [gong2025VEMthermomechanical]. VEM accommodates arbitrary polygonal elements that simplify complex geometry discretization and enables local refinement through non-matching meshes without compromising overall mesh quality. This capability reduces element count and computational cost while preserving solution accuracy, making VEM particularly suitable for multi-scale problems where traditional meshing approaches become prohibitively expensive. Non-matching meshes refer to the situation where, during the independent discretization of two or more adjacent regions within a computational domain, the positions of interface nodes between these regions do not align, leading to misaligned nodes and edges along their shared interface. During the assembly process, nodes from one region’s interface are inserted into the interface elements of the adjacent region based on their coordinates, and any duplicate nodes are merged. Consequently, the interface nodes become shared between both regions’ interface elements, thereby enforcing the continuity of physical quantities such as displacements at the interface.
@@
The VEM was first introduced by Beirao da Veiga et al. [beirao2013basic] for Poisson’s equation and has since been expanded across multiple engineering applications. Early validation studies by Artioli et al. [artioli2017arbitrary] demonstrated VEM’s accuracy through computational benchmarks including Cook’s membrane tests. Subsequent research [sorgente2022role, sorgente2022polyhedral, van2020virtual, van2021virtual] has confirmed VEM’s geometric robustness and convergence properties for arbitrary element shapes, including non-convex configurations. VEM applications have extended beyond linear elasticity to encompass hyperelastic materials [xu2024high, chi2017some], contact mechanics [shen20222, cihan2022virtual], elastoplasticity [hudobivnik2019low, wriggers2017low], and phase field fracture modeling [liu2023virtual]. Dhanush et al. [dhanush2019implementation] pioneered VEM implementation for thermomechanical coupling using Abaqus UEL framework. Recently, Gong et al. [gong2025VEMthermomechanical] analyzed thermoelastic problems using VEM, presenting both the standard virtual element method and a stabilization-free virtual element method [XU2024116826] for thermomechanical behavior analysis in electronic packaging structures with geometric multi-scale features. Their work demonstrated that VEM’s polygonal mesh flexibility enables localized mesh modifications without affecting global mesh structure, making it particularly effective for electronic packaging reliability analysis involving complex geometries. However, VEM requires projection operators and stabilization terms that increase algorithmic complexity and may result in higher local computational costs compared to conventional FEM.
@@ 2.1 Governing equations
$\nabla\cdot(\lambda\nabla T)+Q=0,$ (3)
where $\lambda$ is the thermal conductivity and $T$ is the temperature field, $Q$ is the internal heat generation per unit volume. The boundary conditions on $\partial\Omega=\partial\Omega_{T}\cup\partial\Omega_{q}$ are expressed as
@@
where $\mathbf{I}$ is the identity tensor, $\alpha$ is the coefficient of thermal expansion, and $T_{0}$ is the reference (initial) temperature at which thermal strains are zero.
@@ 2.2 Weak formulation
$\lambda_{k}a_{T}(T,\varphi)+\int_{\Omega}Q\varphi\mathrm{~d}\Omega=\ell_{T}(\varphi),$ (12)
where $\varphi$ is the test function (virtual temperature variation). For simplicity, we assume $Q=0$ (no internal heat generation) in this study.
The corresponding bilinear and linear forms are defined as
@@ 3.1.1 Thermal analysis
For the finite element discretization of the thermal problem, we employ the weak formulation established in Sec. $2.$ The domain is discretized into finite elements such that $\Omega\equiv\text{int}\left(\bigcup_{i}E_{i}\right)$, where the operator ‘int’ denotes the interior of a set, $\Omega$ represents the union of the interior regions of all elements $E_{i}$, excluding the boundaries. And element Ei is chosen as a four-node quadrilateral isoparametric element.
@@ 3.1.2 Thermoelastic analysis
For the finite element discretization of the mechanical problem, we employ the weak formulation established in Sec. $2.$ Within each element, the displacement field is interpolated using the same shape functions as the thermal analysis
@@ 3.2.1 Thermal analysis
In Eq. (41), the first term on the right-hand side represents the consistency term computed from the integral of projected functions in element $E$. The second term is the stabilization term that represents the energy contribution of projection residuals $T^{h}-\Pi^{\nabla}T^{h}$ and $\varphi^{h}-\Pi^{\nabla}\varphi^{h}$ . This stabilization term is essential in VEM formulations to ensure that the stiffness matrix is positive definite and full rank.
@@
$\mathbf{K}_{E}^{th,\mathrm{s}}=\tau^{h}\,\mathrm{tr}\left(\mathbf{K}_{E}^{th,\mathrm{c}}\right)(\mathbf{I}-\bm{\Pi})^{\mathrm{T}}(\mathbf{I}-\bm{\Pi}),$ (44)
@@ 3.2.2 Thermoelastic analysis
where $\mathbf{v}^{h}$ represents the discrete vector function within the element and $\Delta$ represents the Laplace Operator such that $\Delta\mathbf{v}^{h}=\nabla^{2}\mathbf{v}^{h}=\sum_{i=1}^{n}\frac{\partial^{2}\mathbf{v}^{h}}{\partial x_{i}^{2}}$.
@@
where $\tilde{\bm{\Pi}}^{*\nabla}$ is the Ritz matrix representation of the projection operator $\tilde{\bm{\Pi}}^{\nabla}$ can be solved from Eq. (54[xu2024high]. The thermal strain $\hat{\bm{\varepsilon}}_{\mathrm{th}}=\alpha(T-T_{0})[1,1,0]^{\mathrm{T}}$ is expressed in Voigt form.
@@ 4.3.2 Thermal analysis phase
The thermal analysis phase implements the heat conduction formulation established in Sec. $2.$ For the FE domain, standard shape functions are constructed and element thermal stiffness matrices $\mathbf{K}_{\mathrm{FE}}^{th,E}$ are computed using the procedures described in Sec. 3.1.1. Simultaneously, the VE domain employs projection operators to calculate element stiffness matrices $\mathbf{K}_{\mathrm{VE}}^{th,E}$ following the VEM formulation.
@@ 4.3.3 Thermoelastic analysis phase
The mechanical analysis phase utilizes the computed temperature field as thermal loading input, implementing the thermomechanical coupling described in Sec. $2.$ Following procedures analogous to the thermal analysis, shape functions are constructed and element mechanical stiffness matrices are computed for both domains: $\mathbf{K}_{\mathrm{FE}}^{u,E}$ for the finite element region and $\mathbf{K}_{\mathrm{VE}}^{u,E}$ for the virtual element region.
@@
For the electronic packaging structures considered in this work, material properties do not exhibit significant temperature dependence within the operating range, allowing us to employ this decoupled approach without loss of accuracy. By solving the thermal and mechanical problems sequentially, we can utilize specialized solvers optimized for each individual field, which enhances numerical stability and computational efficiency compared to solving the fully coupled system.
@@ 5.1 Square plate with circular hole
Three different numerical approaches were employed to solve this problem: VEM, FEM, and the proposed FE-VE coupling method. The domain is partitioned such that the regular regions employ finite elements while the area near the circular hole utilizes virtual elements to handle the geometric complexity. The resulting von Mises stress distribution contours are presented in Fig. 5. The stress contours demonstrate excellent agreement among all three methods, particularly around the critical stress concentration region near the circular hole. Fig. 5a shows the VEM solution with its characteristic polygonal mesh that naturally adapts to the curved boundary geometry. Fig. 5b presents the FEM solution employing a structured quadrilateral mesh, while Fig. 5c illustrates the FE-VE coupling approach that effectively combines the advantages of both methods. The stress magnitude distributions and overall stress patterns are highly consistent across all three approaches, confirming the accuracy and reliability of the proposed coupling methodology.
@@
The stress distributions along the line $L_{1}$ (given in Fig. 4) were extracted for FEM, VEM, and FE-VE method under different Dofs, as illustrated in Fig. 6. The results demonstrate excellent agreement among all three approaches, confirming the accuracy and consistency of the proposed coupling methodology. Additionally, the computational times for the three methods were compared, as presented in Table 2. From Table 2, it is evident that computational time increases with the number of degrees of freedom (nDofs) for all three methods, which is expected behavior. When comparing methods with similar nDof (approximately 10,700), FEM outperforms VEM with computational times of 0.61 seconds and 0.78 seconds, respectively. This performance gap arises primarily from the complex internal projection operations and stabilization terms required by VEM, which lead to significantly higher computational cost per element. Notably, the coupled FE-VE method achieves the highest efficiency, requiring only 0.47 seconds. By leveraging VEM’s ability to handle complex geometries with fewer elements and FEM’s computational efficiency in regular regions, this hybrid strategy effectively circumvents the high cost of pure VEM in complex domains and the need for excessive mesh refinement with pure FEM in simple regions, resulting in superior overall performance.
@@
where NN represents the total number of nodes in the calculation, $\bm{\sigma}_{\text{num}}$ denotes the numerical solution of nodal stress, and $\bm{\sigma}_{\text{ref}}$ represents the reference solution of nodal stress obtained from mesh refinement (nDof =334,254) in Abaqus.
@@
Fig. 7 presents the MRE convergence comparison for FEM, VEM, and FE-VE coupling method as the nDof increases from approximately 5,000 to 25,000. The convergence analysis reveals distinct performance characteristics for each method, with convergence rates indicated by the slope annotations in the figure. The convergence rates indicate that the FE-VE coupling method actually achieves slightly better convergence behavior than pure FEM, while VEM shows the most gradual convergence. All three methods demonstrate stable convergence trends, confirming the reliability of the proposed coupling strategy. The FE-VE approach maintains good accuracy while providing the additional advantage of geometric flexibility for complex domain geometries.
@@ 5.2 Thermoelastic analysis of thick-walled cylindrical structure
We analyze a homogeneous thick-walled cylinder under thermal loading to validate the FE-VE coupling method. The analytical solution for this problem serves as a benchmark for accuracy assessment. Fig. 8 illustrates the geometry and boundary conditions of this verification case.
@@
Figure 8: Thick-walled cylinder configuration: (a) Full geometry with inner radius $r_{a}$, outer radius $r_{b}$, and applied temperatures $T_{a}$ and $T_{b}$; (b) Mesh employed in the FE-VE coupled method
@@
where NN is the number of evaluation points, and $X_{e}$ and $X$ denote the analytical and numerical solutions, respectively.
@@
Fig. 9 presents the convergence analysis comparing FE-VE coupling, pure VEM, and pure FEM methods, showing RMS errors for temperature as functions of degrees of freedom. The convergence rates, indicated by the slope annotations in the figure, demonstrate that the VEM achieves a convergence rate of 1.02, while the FE-VE coupling method exhibits a rate of 1.01, both outperforming pure FEM with a rate of 0.92. The proposed FE-VE coupling method maintains nearly optimal convergence characteristics comparable to pure VEM, while providing the additional advantage of geometric flexibility through strategic domain partitioning. All three methods show consistent error reduction with mesh refinement, confirming the stability and reliability of the coupling approach.